OpenStudy (anonymous):

Solve the equation. ((x-21)/7)+((x+4)/4)=x-4

OpenStudy (anonymous):

X = 56/17

OpenStudy (anonymous):

i need work..

OpenStudy (anonymous):

4x - 84 +7x + 28 = 28x - 112 17x = 112 - 56 17x = 56 x = 56/17